# onlineReferences

### v5

@@ Abstract
Many studies exploit variation in policy adoption timing across units as an instrument for treatment. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. Moreover, we propose a credible estimation method in this design that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings, estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, the conventional approach to implement DID-IV designs, yields a negative estimate. By contrast, our estimation method indicates a substantial gain from schooling.
@@
[…]
@@ 1   Introduction
To identify the effect of a treatment on an outcome, many studies have exploited variation in policy adoption timing across units as an instrument for treatment. For example, to estimate returns to schooling in Indonesia, Duflo (2001) leverages variation in the timing of introduction of school construction program across regions as an instrument for education attainment. Similarly, to estimate the causal link between parents’ and children’s education attainment, Black et al. (2005) leverage variation in the timing of implementation of school reforms across municipalities as an instrument for parents’ education attainment. Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that the timing variation of a policy shock is used for identifying the treatment effects. Meanwhile, it also differs from DID designs, in that the variation is used to construct an instrument rather than a treatment.
@@
In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV). We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this empirical application, we demonstrate that the estimation method matters in this design in practice.
@@
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument during the two periods (unexposed group), while some units become exposed in the second period (exposed group). [^1] Note that this two-period/two-group ($2\times 2$) setting is already considered in de Chaisemartin (2010) and Hudson et al. (2017), and our $2\times 2$ DID-IV design is mainly based on these studies. In this paper, we revisit $2\times 2$ DID-IV designs for two reasons. First, we aim to complement de Chaisemartin (2010) and Hudson et al. (2017) along two dimensions. Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs, and uncover the additional identifying assumption not give in the previous literature. In addition, we investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in DID-IV designs. Second, we aim to compare $2\times 2$ DID-IV designs to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018). While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) treat the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups. Our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period. We show that in this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the LATET.
In DID-IV designs, the parallel trends assumption in the outcome requires careful assessment because it does not depend on the potential outcomes, but depends on the outcome under no exposure to the instrument, which we call an unexposed outcome. Give this, we next clarify the interpretation of this assumption in DID-IV settings. Specifically, we first point out that, in DID-IV designs, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of a policy shock (instrument). Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers. Subsequently, we show that the average time trend in an unexposed outcome equals a weighted average of the effect of time on an outcome through treatment in each time group. We call this an expected time gain, and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
@@
de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period. Because the target parameter differs from each other, the identifying assumptions in dCDH are also different from those used here.
@@
Motivated by the differences between this article and dCDH, we next examine the detailed connections between DID-IV and Fuzzy DID, and point out the issues inherent in Fuzzy DID. First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units, and these restrictions are asymmetric between unexposed and exposed groups. Second, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers. This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher. This is because the SLATET is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by instrument.
@@
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In most DID-IV applications, researchers exploit variation in policy adoption timing across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), phased-in introduction of head starts across states (e.g. Johnson and Jackson (2019)), or gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions. Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument. Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings, and state the interpretation of each assumption.
@@
We extend our DID-IV designs in several directions. We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment. We also consider extensions to repeated cross sections. Lastly, we consider the case where we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
@@
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable (TWFEIV) regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$. We also discuss the procedure of pre-trends tests to check the validity of the parallel trends assumptions in DID-IV designs.
@@
We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment. In this application, we first assess the plausibility of the DID-IV identification strategy implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. Finally, we use our estimation method to reassess returns to schooling in the UK. We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighting estimate indicates roughly a $20\%$ gain from schooling in the UK.
Overall, this paper provides a new econometric framework by combining the IV techniques with DID designs. Our DID-IV design can be applied to various empirical settings. First, it can cope with the general adoption process of the treatment, which the canonical DID designs cannot, when we are interested in the effect of a treatment on an outcome. For instance, suppose that some units already receive the treatment in period 00 before the policy shock that distributes the treatment in period $1$. In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data. By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET. In other words, we can view that DID-IV is DID with noncompliance of the assigned treatment.
Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs. For instance, suppose that the treatment is education attainment and we are interested in returns to schooling as in Duflo (2001). In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality. By contrast, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument). Once these groups are constructed, one can identify the local average treatment effect by comparing the evolution of the treatment and the outcome between the two groups.
Finally, our DID-IV design allows a researcher to address the endogeneity issue in a given application. In many empirical settings, a researcher often fails to identify the treatment effects due to the endogenous adoption of the treatment or the potential violation of the parallel trends assumption under DID designs. In this situation, our DID-IV design offers an alternative identification strategy, constructing an instrument based on the timing variation of a policy shock. Notably, in this design, the identifying variation comes not from the random assignment of a policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome. This identification strategy would be less demanding than IV designs in certain applications, and enhance the internal and external validity of the IV estimate in practice (Imbens and Angrist (1994); Heckman and Vytlacil (2001); Heckman and Vytlacil (2005)).
@@ 1.1   Related literature
Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018); de Chaisemartin et al. (2024)) and contributes to this literature in three important ways. [^2] While Ye et al. (2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to capture this parameter. We therefore see that Ye et al. (2023) are distinct from the recent DID-IV literature.
@@
First, this paper investigates the detailed connections between DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that if parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption hold, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET). Following de Chaisemartin (2010), Hudson et al. (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings. Building on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs.
@@
In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al. (2017) in two ways. Specifically, the framework we consider is mainly based on de Chaisemartin (2010) and Hudson et al. (2017). Nevertheless, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover the additional identifying assumption not given in the previous literature. We also investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID, and point out the issues inherent in Fuzzy DID. The discussion organizes the relationship between the two designs, which essentially treat the same setting, in the previous literature, and indicates the relative advantages of DID-IV over Fuzzy DID.
@@
Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design. In practice, empirical researchers often leverage variation in policy adoption timing across units as an instrument for treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). No previous study, however, has extended $2\times 2$ DID-IV designs to such important settings. In this paper, we formalize the underlying identification strategy as staggered DID-IV designs, and establish the target parameter and the identifying assumptions. Our staggered DID-IV designs allow the practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
@@
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or the treatment evolves over time. [^3] See also de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator of the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally in Section 3.4 of their Web Appendix. Our proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
@@
In work related to this paper, de Chaisemartin et al. (2024) also consider DID-IV designs following de Chaisemartin (2010) and Hudson et al. (2017). This paper is distinct from theirs in three ways. First, they consider continuous treatment with continuous but static instrument, while we consider binary or ordered treatment with binary but dynamic instrument. Second, they consider multiple period settings with potentially non-staggered instrument, while we consider multiple periods settings with staggered instrument. Third, they consider the Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
@@
Our paper is also related to a growing DID literature in two ways. First, this paper provides an alternative identification strategy to DID designs, addressing cases where the parallel trends assumption is not satisfied or no control group exists in a given application. When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often concern that the parallel trends assumption is not plausible in practice. To address this, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g. Miller and Segal (2019)). In this paper, we formalize the underlying identification strategy as instrumented difference-in-differences. Our DID-IV design is a natural extension of DID design, in that the timing variation is used to construct an instrument rather than a treatment.
@@
Second, in line with the DID literature, this paper develops a reliable estimation method in DID-IV designs in the presence of heterogeneous treatment effects. Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects. (Athey and Imbens (2022); Borusyak et al. (2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)). Some of them have proposed a credible estimation method that delivers a sensible estimand and is robust to treatment heterogeneity. Along with the recent DID literature, this paper proposes an alternative estimation method to the TWFEIV regression, and illustrate its usefulness through an empirical application.

### v6

@@ Abstract
Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. Moreover, we propose a credible estimation method in this design that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings, estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, the conventional approach to implement DID-IV designs, yields a negative estimate. By contrast, our estimation method indicates a substantial gain from schooling.
@@
In this online appendix, we first present additional identification results. Second, we derive the influence function of our STS estimator in repeated cross section settings. Third, we decompose the TWFEIV estimand in the setting of Oreopoulos (2006). Finally, we investigate the connections between DID-IV and Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018).
@@ 1  Introduction
To identify the effect of a treatment on an outcome, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment. For example, to estimate returns to schooling in Indonesia, Duflo (2001) exploits variation in the timing of the introduction of a school construction program across regions as an instrument for educational attainment. Similarly, to estimate the causal link between parents’ and children’s educational attainment, Black et al. (2005) exploit variation in the timing of the implementation of school reforms across municipalities as an instrument for parents’ educational attainment. Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that variation in the timing of a policy shock is used to identify treatment effects. At the same time, it differs from DID designs in that this variation is used to construct an instrument rather than a treatment.
@@
In this paper, we formalize this identification strategy as an instrumented difference-in-differences (DID-IV). We define the target parameter and the identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this application, we show that the choice of estimation method matters in practice.
@@
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument in either period (the unexposed group), while others become exposed in the second period (the exposed group). [^1] Note that this two-period/two-group ($2\times 2$) setting has already been considered in de Chaisemartin (2010) and Hudson et al. (2017), and our $2\times 2$ DID-IV design builds on these studies. In this paper, we revisit $2\times 2$ DID-IV designs for two reasons. First, we aim to complement de Chaisemartin (2010) and Hudson et al. (2017). Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover an additional identifying assumption that is not made explicit in the previous literature. Second, we compare $2\times 2$ DID-IV designs to the Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018). While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) study the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this setting, our DID-IV design relies on a monotonicity assumption and parallel trends assumptions for both the treatment and the outcome between the exposed and unexposed groups. Our target parameter is the local average treatment effect on the treated (LATET), which measures the treatment effect for units that belong to the exposed group and are induced to receive the treatment by the instrument in the second period. We show that, in this design, a Wald-DID estimand—defined as the ratio of the DID estimand of the outcome to that of the treatment—identifies the LATET.
@@
de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this paper and dCDH lies in the definition of the target parameter. While we focus on the LATET, dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period. Because the target parameters differ, the identifying assumptions adopted in dCDH also differ from those used in this paper.
@@
Motivated by these differences, we next examine the detailed connections between DID-IV and Fuzzy DID and discuss their implications for treatment adoption behavior and the interpretation of the target parameter. We first show that the identifying assumptions under Fuzzy DID impose stronger restrictions on treatment adoption behavior than those under DID-IV. Under these restrictions, we then demonstrate that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two distinct causal parameters. One parameter captures the treatment effects among the subpopulation of the compliers in the sense of Imbens and Angrist (1994), while the other captures the treatment effects for time compliers—units whose treatment status is affected by time but not by the instrument. This decomposition result has an important implication: even when the instrument is directly linked to the policy change of interest, the SLATET may fail to be a policy-relevant parameter (Heckman and Vytlacil (2001)).
@@
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for example, from the staggered adoption of school reforms across municipalities or countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), the phased-in introduction of Head Start across states (e.g. Johnson and Jackson (2019)), or the gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We refer to this identification strategy as a staggered DID-IV design, and establish the corresponding target parameter and identifying assumptions. Specifically, we first partition units into mutually exclusive and exhaustive cohorts based on the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT). This parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs and measures the treatment effects for units that belong to cohort $e$ and are the compliers at a given relative period $l$ following the initial exposure to the instrument. Finally, we introduce two Wald–DID estimands that use either never-exposed cohorts or not-yet-exposed cohorts as control groups. For each estimand, we extend the identification assumptions from the $2\times 2$ DID–IV setting to multiple time periods, and show that, under these assumptions, the corresponding Wald–DID estimand identifies the CLATT at each relative period following the initial exposure to the instrument.
@@
We extend our DID-IV framework along two dimensions. First, we show that it can be applied to settings with a non-binary, ordered treatment. Second, we consider extensions to repeated cross sections.
@@
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they typically implement it using two-way fixed effects instrumental variable (TWFEIV) regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In a companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize treatment effects if the effect of the instrument on the treatment or on the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following Callaway and Sant’Anna (2021), we propose a weighting scheme to summarize the treatment effects, where the weight reflects the share of compliers in a given relative period $l$ in cohort $e$. We also discuss the procedure of pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
@@
We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of the implementation of school reforms between Britain and Northern Ireland as an instrument for education attainment. In this application, we first assess the plausibility of the DID-IV identification strategy implicitly adopted by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. Finally, we use our estimation method to reassess returns to schooling in the UK. We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighted estimate indicates roughly a $20\%$ gain from schooling in the UK.
@@
Our paper is related to the recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018); de Chaisemartin et al. (2024); Chen et al. (2025); Helmers et al. (2025)) and contributes to this literature in three ways. [^2] While Ye et al. (2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they impose strong assumptions for the Wald-DID estimand to identify this parameter. We therefore view Ye et al. (2023) as distinct from the recent DID-IV literature.
@@
First, this paper investigates the detailed connections between DID-IV and the Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that, under parallel trends assumptions for both the treatment and the outcome, together with a monotonicity assumption, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET). [^3] Blundell and Dias (2009) also consider a DID–IV setting with a binary treatment and a binary instrument in Subsection F.2 of Section III. There are two differences between de Chaisemartin (2010) and Blundell and Dias (2009). First, while de Chaisemartin (2010) focus on the LATET as the target parameter, Blundell and Dias (2009) focus on the switcher local average treatment effect (SLATET), as in dCDH. Second, while de Chaisemartin (2010) impose a parallel trends assumption on unexposed outcomes, Blundell and Dias (2009) impose a parallel trends assumption on untreated outcomes. Following de Chaisemartin (2010), Hudson et al. (2017) also study $2\times 2$ DID-IV designs with a non-binary, ordered treatment. [^4] Recently, Chen et al. (2025) extend our DID-IV framework by allowing for covariates. Helmers et al. (2025) consider an extension to the continuous treatment case. Building on de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
@@
In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al. (2017). Specifically, while the framework we consider here is mainly based on de Chaisemartin (2010) and Hudson et al. (2017), we explicitly introduce the instrument path into $2\times 2$ DID-IV designs and uncover an additional identifying assumption not given in the previous literature.
Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID. Specifically, we clarify the differences between DID-IV and Fuzzy DID and discuss their implications for treatment adoption behavior and the interpretation of the target parameter.
@@
Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, which we refer to as staggered DID-IV designs. In practice, empirical researchers often leverage variation in policy adoption timing across units as an instrument for treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). However, no previous study has extended $2\times 2$ DID-IV designs to such important settings. In this paper, we formalize the underlying identification strategy as staggered DID-IV designs, and establish the target parameter and the identifying assumptions. Our staggered DID-IV designs allow practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
@@
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV designs in practice, they commonly implement this design via TWFEIV regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In a companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or on the treatment evolves over time. [^5] See also de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator of the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally in Section 3.4 of their Web Appendix. Our proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
@@
In work related to this paper, de Chaisemartin et al. (2024) also study DID-IV designs building on de Chaisemartin (2010) and Hudson et al. (2017). Our paper is distinct from theirs in three ways. First, they consider continuous treatment with continuous but static instrument, while we consider binary or ordered treatment with binary but dynamic instrument. Second, they consider multiple period settings with potentially non-staggered instrument, while we consider multiple periods settings with staggered instrument. Third, they view Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
@@
Our paper is also related to the recent DID literature in two ways. First, this paper provides an alternative identification strategy to DID designs, addressing settings in which the parallel trends assumption is unlikely to hold or no credible control group exists. When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often worry that the parallel trends assumption is implausible in practice. To address this concern, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g., Miller and Segal (2019)). In this paper, we formalize the underlying identification strategy as instrumented difference-in-differences. Our DID-IV design can be viewed as a natural extension of DID designs, in that the timing variation is used to construct an instrument rather than a treatment.
@@
Second, in line with the DID literature, this paper develops a reliable estimation method for DID-IV designs in the presence of heterogeneous treatment effects. Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al. (2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021); Callaway et al. (2024)). Some of these studies propose credible estimation methods that deliver a sensible estimand and are robust to treatment heterogeneity. In the same spirit as the recent DID literature, this paper proposes an alternative to the TWFEIV regression and illustrates its usefulness through an empirical application.